40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Wed Sep 21 17:01:59 2022, Paco, Update, BHD, BH55 RFPD installed - part I IMG_3760.jpegIMG_3764.jpeg
    Reply  Thu Sep 22 19:51:58 2022, Anchal, Update, BHD, BH55 LSC Model Updates - part II 
       Reply  Fri Sep 23 19:04:12 2022, Anchal, Update, BHD, BH55 LSC Model Updates - part III 
          Reply  Thu Sep 29 18:01:14 2022, Anchal, Update, BHD, BH55 LSC Model Updates - part IV 
             Reply  Fri Sep 30 18:30:12 2022, Anchal, Update, ASS, Model Changes 
                Reply  Wed Oct 12 14:01:23 2022, Radhika, Update, ASS, Model Changes YARM_ASS.png
                   Reply  Wed Oct 12 18:18:16 2022, rana, Update, ASS, Model Changes 
    Reply  Fri Sep 23 14:10:19 2022, Radhika, Update, BHD, BH55 RFPD installed - part I IMG_3766.jpegIMG_3770.jpegIMG_3773.jpeg
       Reply  Fri Sep 23 18:31:46 2022, rana, Update, BHD, BH55 RFPD installed - part I 
          Reply  Mon Sep 26 11:39:37 2022, Paco, Update, BHD, BH55 RFPD installed - part II PXL_20220926_175010061.MOTION.jpgBH55_Transimpedance_Measurement.pdf
             Reply  Thu Oct 6 11:12:14 2022, Anchal, Update, BHD, BH55 RFPD installation complete 
Message ID: 17152     Entry time: Thu Sep 22 19:51:58 2022     In reply to: 17150     Reply to this: 17157
Author: Anchal 
Type: Update 
Category: BHD 
Subject: BH55 LSC Model Updates - part II 

I updated follwoing in teh rtcds models and medm screens:

  • c1lsc
    • Added reading of ADC0_20 and ADC0_21 as demodulated BHD output at 55 MHz, I and Q channels.
    • Connected BH55_I and BH55_Q to phase rotation and creation of output channels.
    • Replaced POP55 with BH55 in the RFPD input matrix.
    • Send BH55_I and BH55_Q over IPC to c1hpc
    • Added BH55 RFPD model in LSC screen, in RFPD input matrix, whitening box. Some work is still remaining.
  • c1hpc
    • Added recieving BH55_I and BH55_Q.
    • Added BH55_I and BH55_Q to sensing matrix through filter modules. Now these can be used to control LO phase.
    • Added BH55 signals to the medm screen.
  • c1scy
    • Updated SUS model to new sus model that takes care of data acquisition rates and also adds BIASPOS, BIASPIT and BIASYAW filter modules at alignment sliders.

Current state:

  • All models built and installed without any issue or error.
  • On restarting all models, I first noticed 0x2000 error on c1lsc, c1scy and c1hpc. But these errors went away with doing daqd restart on fb1.
  • BH55 FM buttons are not connected to antialiasing analog filter. Need to do this and update medm screen accordingly.
  • The IPC from c1lsc to c1hpc is not working. One sender side, it does not show any signal which needs to be resolved.
ELOG V3.1.3-